# TimeDiT: General-purpose Diffusion Transformers for Time Series Foundation Model

### v1

@@ 5 Experiments
To comprehensively assess our time series foundation model, we evaluate on a diverse set of tasks that reflect real-world challenges and applications. We begin by testing the model’s performance in practical scenarios and its ability to integrate domain knowledge. This includes handling missing data and multi-resolution forecasting on customized datasets, which allows us to evaluate the model’s robustness in situations that frequently occur in real-world applications, as well as physics-informed modeling crucial for scientific and engineering domains(Yuan and Qiao, 2024), which uses 6 practical partial differential equations (PDEs). We then assess the model’s capabilities in well-established benchmarking tasks across various fields such as finance, healthcare, and industrial monitoring. These tasks include forecasting on Solar, Electricity, Traffic, Taxi, and Exchange datasets(Tashiro et al., 2021; Rasul et al., 2021) to evaluate temporal dependency modeling, imputation on ETTh, ETTm, Weather and Electricity datasets(Zhou et al., 2021) to assess the handling of missing data, anomaly detection on MSL, SMAP, SWaT, SMD, and PSM datasets(Xu et al., 2021; Zhao et al., 2020) to gauge sensitivity to unusual patterns, and synthetic data generation on Stock, Air Quality, and Energy datasets(Yoon et al., 2019; Desai et al., 2021) to test understanding of underlying distributions. By evaluating these diverse tasks, we can demonstrate that our model truly serves as a foundation for various time series applications, potentially reducing the need for task-specific models.
Figure 2: Visualization of miss value (a) and multi resolution (b) forecasting results on the Exchange dataset. Compared between our model TimeDiT and state-of-the-art diffusion-based methods. The x-axis number in (b) is the number of different sampling resolutions in the multivariate input.
5.1 Practical Scenarios: Missing Data and Multi-Resolution Forecasting
To simulate more realistic scenarios in time series tasks, we introduced two additional challenges: missing values and multi-resolution data. These conditions are common in real-world applications and test a model’s robustness and adaptability. For the missing value scenario, we created datasets with various missing ratios, simulating incomplete data often encountered in practice. In the multi-resolution setting, we sampled each individual time series within the multivariate dataset at different resolutions, reflecting the diverse sampling frequencies often present in real-world data collection.
Figure  2 illustrates TimeDiT’s performance in realistic scenarios, showcasing its effectiveness across different sampling frequencies on the Exchange dataset. In Figure  2 (a), we observe TimeDiT’s superior performance in handling missing data. As the missing ratio increases from 5% to 50%, TimeDiT maintains the lowest CRPSsum across all scenarios, indicating its robustness to data gaps. The performance gap between TimeDiT and other models widens as the missing ratio increases, highlighting its effectiveness in more challenging conditions. Figure  2 (b) demonstrates TimeDiT’s ability to manage multi-resolution data, where it maintains a clear performance advantage as the number of different sampling resolutions increases from 2 to 6. This demonstrates its ability to effectively integrate and forecast time series data sampled at varying frequencies. These findings underscore TimeDiT’s potential as a practical and versatile tool for time series forecasting in diverse, challenging scenarios that more closely resemble real-world applications.
@@ 5.1 Practical Scenarios: Missing Data and Multi-Resolution Forecasting
Table 1: Physics informed results: This table presents results of PDE forecasting for various sampling strategies, demonstrating the model’s ability to incorporate physics constraints without updating the foundation model. Metrics shown include point estimates metrics, which is MSE, RMSE and MAE; probabilistic metrics, CRPS, for uncertainty quantification. Lower values indicate better performance and closer adherence to physical laws.
5.2 Domain Knowledge Integration: Physics-Informed TimeDiT
A key advantage of our approach is the ability to directly incorporate physics knowledge into the pretrained foundation model without additional fine-tuning. This is possible because the diffusion model reconstructs the entire process, allowing for seamless integration of PDE-based constraints during inference. By encoding the known physical laws governing the system into the sampling process, we can guide the model towards more physically consistent and accurate predictions. In this section, we evaluate how effectively our pre-trained foundation model can integrate physics-informed knowledge into time series forecasting without the need for fine-tuning. We study four 1D partial differential equations (PDEs) from  (Takamoto et al., 2022): general Navier-Stokes Equations, Kolmogorov Flow (a specific case of Navier-Stokes Equations), Advection Equations, Burgers Equations, Diffusion Soeption and Computational Fluid Dynamics (CFD). These equations are used to generate synthetic data with random initial conditions, and we apply diffusion models to forecast time series based on data from a historical window.
@@ 5.2 Domain Knowledge Integration: Physics-Informed TimeDiT
Table 1 presents the results, including both mean error and error bars. The table clearly demonstrates that our proposed model editing solution, which incorporates physics knowledge, significantly outperforms previous sampling strategies such as DDPM (Ho et al., 2020), DDIM (Song et al., ), and TS Diffusion’s Self-Guidance (Kollovieh et al., 2023). By leveraging domain-specific physical information, our approach achieves substantial performance improvements over these baselines, highlighting the effectiveness of integrating physics-informed priors into the diffusion model sampling process. This performance gain underscores the potential of combining pretrained foundation models with domain-specific knowledge. Our method offers a flexible framework for enhancing time series forecasting in scientific and engineering applications where the underlying physical laws are partially known. It demonstrates that by bridging the gap between data-driven approaches and physics-based modeling, we can achieve more accurate and physically consistent predictions without the computational overhead of retraining or fine-tuning the entire model. The ability to easily incorporate physics knowledge into a pretrained foundation model represents a significant advance in the field of scientific machine learning. It opens up new possibilities for rapid adaptation to specific physical systems and phenomena, potentially accelerating research and discovery in various scientific domains.
5.3 Forecasting on Full-shot Setting and Zero-shot Setting
In the forecasting task, we conduct two types of experiments. First, we compare our proposed TimeDiT with baselines in a full-shot setting, where models are trained and tested on separate datasets. This approach evaluates their performance on conventional time series forecasting tasks, ensuring that models can effectively learn and generalize from complete data. Second, we assess TimeDiT as a foundation model in a zero-shot setting, comparing it to previous transformer-based time series models. This setting is crucial as it tests the model’s ability to generalize and adapt to entirely new datasets without prior exposure, highlighting its robustness and versatility. Together, these experiments provide a holistic view of TimeDiT’s capabilities, addressing both specialized performance and broad applicability in time series forecasting.
Solar Electricity Traffic Taxi Exchange
Deterministic DLinear 0.432(0.002) 0.033(0.000) 0.070(0.001) 0.177(0.000) 0.011(0.001)
PatchTST 0.457(0.019) 0.037(0.002) 0.405(0.001) 0.190(0.005) 0.026(0.001)
Latent ODE 0.445(0.002) 0.140(0.017) 0.095(0.004) 0.181(0.006) 0.013(0.001)
GPT2(6) 0.467(0.002) 0.033(0.001) 0.069(0.001) 0.187(0.001) 0.013(0.001)
Probabilistic GP-copula 0.337(0.024) 0.024(0.002) 0.078(0.002) 0.208(0.183) 0.007(0.000)
TransMAF 0.301(0.014) 0.021(0.011) 0.056(0.001) 0.179(0.002) 0.005(0.003)
TimeGrad 0.287(0.020) 0.021(0.001) 0.044(0.006) 0.114(0.020) 0.006(0.001)
CSDI 0.298(0.004) 0.017(0.000) 0.020(0.001) 0.123(0.003) 0.007(0.001)
Diffusion-TS 0.286(0.003) 0.019(0.002) 0.097(0.001) 0.303(0.004) 0.009(0.001)
TimeDiT 0.278(0.001) 0.017(0.000) 0.019(0.000) 0.123(0.001) 0.005(0.001)
Table 2: Forecasting results on CRPSsum with full shot setting.
Solar Electricity Traffic Taxi Exchange
Zero-shot TEMPO 0.581(0.002) 0.081(0.003) 0.147(0.000) 0.400(0.001) 0.030(0.001)
Moirai small 0.884(0.005) 0.079(0.002) 0.215(0.000) 0.463(0.001) 0.007(0.000)
Moirai base 0.948(0.002) 0.072(0.002) 0.191(0.001) 0.428(0.000) 0.012(0.000)
Moirai large 1.042(0.002) 0.039(0.001) 0.111(0.000) 0.597(0.000) 0.011(0.000)
LagLLama 0.690(0.005) 0.065(0.005) 0.275(0.001) 0.620(0.003) 0.024(0.001)
TimeDiT 0.457(0.002) 0.026(0.001) 0.185(0.010) 0.398(0.001) 0.021(0.002)
Table 3: Forecasting results on CRPSsum. Zero-shot implies that the model has not encountered any samples from the above datasets during training.
Table  2 presents the full-shot forecasting results, comparing TimeDiT with state-of-the-art models in two categories: deterministic forecasting models, which are trained with the Student’s t-distribution head to support probabilistic results, and inherently probabilistic time series forecasting models, including diffusion-based models, such as CSDI and non-diffusion-based, such as GP-copula. Our model achieves the lowest CRPSsum on four datasets and the second-best performance on the Taxi dataset. In the zero-shot setting, TimeDiT is compared with the open-sourced foundation models including TEMPO (Cao et al., 2023b), which is pre-trained with Student’s t-distribution head to support probabilistic results, Moirai (Woo et al., 2024b) and LagLLama (Rasul et al., 2023) in Table 3. TimeDiT’s ability to outperform other open-source foundation models in most cases is particularly noteworthy, as it suggests that TimeDiT can be effectively applied to a wide range of time series forecasting tasks across different domains with minimal adaptation.
Datasets ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity 1st Pl
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE Count
DLinear 0.201 0.306 0.142 0.259 0.093 0.206 0.096 0.208 0.052 0.110 0.132 0.260 0
LightTS 0.284 0.373 0.119 0.250 0.104 0.218 0.046 0.151 0.055 0.117 0.131 0.262 0
ETSformer 0.202 0.329 0.367 0.436 0.120 0.253 0.208 0.327 0.076 0.171 0.214 0.339 0
FEDformer 0.117 0.246 0.163 0.279 0.062 0.177 0.101 0.215 0.099 0.203 0.130 0.259 0
Autoformer 0.103 0.214 0.055 0.156 0.051 0.150 0.029 0.105 0.031 0.057 0.101 0.225 0
PatchTST 0.115 0.224 0.065 0.163 0.047 0.140 0.029 0.102 0.034 0.055 0.072 0.183 0
TimesNet 0.078 0.187 0.049 0.146 0.027 0.107 0.022 0.088 0.030 0.054 0.092 0.210 1
GPT2(3) 0.069 0.173 0.048 0.141 0.028 0.105 0.021 0.084 0.031 0.056 0.090 0.207 1
TimeDiT 0.042 0.135 0.042 0.139 0.023 0.098 0.024 0.083 0.031 0.036 0.069 0.174 10
Table 4: Imputation result on 96-length multivariate time series averaged over the four mask ratios. We calculate MSE and MAE for each dataset. Bold indicates best result, Underline indicates the second best result.
5.4 Imputation Task
We conduct experiments on six benchmark time-series datasets: ETTh1, ETTh2, ETTm1, ETTm2, Electtricity, and Weather. We use random mask ratios $\{12.5\%,25\%,37.5\%,50\%\}$ following previous studies’ settings with sequence length set to 96. We finetune on model checkpoints pretrained on solar, traffic, exchange, taxi, Huawei cloud, air quality, and weather (different from the evaluation weather data). Table 4 shows the imputation result averaged over the four mask ratios. TimeDiT achieves the best performance on most dataset. obtaining 10 first places out of the 12 evaluations while the remaining baselines obtained 2 first place count in total. In particular, TimeDiT achived a 39% reduction in MSE and 22% reduction in MAE compared to the strongest baseline on ETTh1 dataset. For full result on each mask ratio, please refer to section F.
5.5 Anomaly Detection Task
We conduct experiments on five real-world datasets from industrial applications: MSL, SMAP, SWaT, SMD, and PSM. The diffusion model, renowned for its proficiency in distribution learning, may inadvertently overfit by reconstructing anomalies alongside normal data points. To counteract this, we opted to bypass pretraining and introduced spectral residue (SR) transformation at the preprocessing stage of TimeDiT . This transformation helps to conceal points most likely to be anomalies and their immediate neighbors. The number of neighbors affected is controlled by the hyperparameter $n_{neighbor}$. The SR method utilizes Fourier Transformation to convert the original time series into a saliency map, thereby amplifying abnormal points, as detailed in (Ren et al., 2019; Zhao et al., 2020). For additional information about this transformation, please see section G. Consistent with prior methodologies, we set the sequence length to be 100 identify anomalies using the 99th percentile of reconstruction errors. During evaluations, we apply standard anomaly adjustments as suggested by (Xu et al., 2018). As demonstrated in table 11, TimeDiT outperforms baseline models on four of the five datasets. In particular, TimeDiT 23.03 points of improvement in terms of F1 score on the SMAP dataset compared to the strongest baseline.
Methods TimeDiT GPT2(6) TimesNet PatchTS. ETS. FED. LightTS DLinear Auto. Anomaly.*
MSL 89.33 82.45 81.84 78.70 85.03 78.57 78.95 84.88 79.05 83.31
SMAP 95.91 72.88 69.39 68.82 69.50 70.76 69.21 69.26 71.12 71.18
SWaT 96.46 94.23 93.02 85.72 84.91 93.19 93.33 87.52 92.74 83.10
SMD 83.28 86.89 84.61 84.62 83.13 85.08 82.53 77.10 85.11 85.49
PSM 97.57 97.13 97.34 96.08 91.76 97.23 97.15 93.55 93.29 79.40
1st Pl Count 4 1 0 0 0 0 0 0 0 0
    We replace the joint criterion in Anomaly Transformer with reconstruction error for fair comparison.
@@ 5.5 Anomaly Detection Task
Table 5: Anomaly Detection result on 100-length multivariate time series. We calculate F1 score as % for each dataset. ’.’ notation in model name stands for transformer. Bold indicates best result, Underline indicates the second best result.
@@
Table 6: Synthetic Generation result on 24-length multivariate time series. We calculate discriminative and predictive score according to Yoon et al. (2019) and results are averaged over five runs. Bold indicates the best performance.
5.6 Synthetic Generation Task
Figure 3: PCA Evaluation of Synthetic Time Series from TimeDiT and other baselines on stock dataset.
@@ 5.6 Synthetic Generation Task
We conduct experiments to synthesize multivariate time series and evaluate performance using the discriminative score and predictive score metrics under a "train on synthetic test on real" experimental setup with sequence length set to 24  (Yuan and Qiao, 2024). We finetune on model checkpoints pretrained on etth1, etth2, ettm1, ettm2, electricity, and weather. Table 6 shows the result on synthetic generation where TimeDiT, in general, consistently generates more realistic synthetic samples compared to baselines, even on high-dimensional energy datasets. This demonstrates TimeDiT’s strength in complex time series synthesis. We visualize synthesis performance using PCA and t-SNE in Appendix E. As shown in Figure 3, TimeDiT’s samples markedly overlap the original data distribution better than other methods. Qualitative and quantitative results confirm TimeDiT’s superior ability to model intricate characteristics for realistic time series synthesis, even on multidimensional, complex datasets.

### v2

@@ 4 Experiments
The development of foundation models for time series analysis presents unique challenges at the intersection of universal representation learning and task-specific adaptation. While traditional foundation models in domains like natural language processing demonstrate strong zero-shot capabilities, time series tasks often require nuanced understanding of domain-specific characteristics despite sharing fundamental temporal dependencies. This observation motivated our comprehensive three-perspective experimental framework: First, we assess the model’s foundational capabilities through both zero-shot transfer and fine-tuning approaches on canonical tasks, including forecasting and imputation on ETTh, Solar, Electricity, Traffic, Taxi, Exchange, ETTm, Weather. Second, we investigate domain-adaptive tasks that require prior knowledge of data characteristics—specifically miss-value and multi-resolution forecasting including Air Quality from climate, MIMIC-III and PhysioNet from healthcare, and NASDAQ from finance, anomaly detection on MSL, SMAP, SWaT, SMD, and PSM datasets(Xu et al., 2021; Zhao et al., 2020), and targeted data generation on Stock, Air Quality, and Energy datasets(Yoon et al., 2019; Desai et al., 2021). These experiments evaluate the model’s capability to bridge the gap between universal temporal patterns and domain-specific requirements. Finally, we examine the model’s capacity for knowledge integration through physics-informed constraints by accurately processing complex partial differential equations (PDEs) (Yuan and Qiao, 2024) and multimodal fusion, testing its ability to incorporate external domain knowledge to enhance its foundational capabilities further.
4.1 Zero-Shot and Finetuning Setting
Forecasting on Zero-shot Setting:
We comprehensively evaluate TimeDiT in zero-shot forecasting environments against leading architectures, encompassing both zero-shot and full-shot models. Using popular benchmark datasets for long-term deterministic forecasting that exclude source overlap with pre-training data, TimeDiT demonstrates exceptional performance across varying prediction horizons (96-720 time steps) as shown in Table 1. The model achieves superior Mean Squared Error (MSE) and Mean Absolute Error (MAE) metrics compared to all baselines, particularly excelling on the ETTh1 dataset where it consistently outperforms all comparative models.
Building upon this success, we extended our investigation to probabilistic forecasting, where TimeDiT demonstrates remarkable capabilities. As evidenced in Table 2, zero-shot TimeDiT achieves superior performance compared to models with fewer parameters in equivalent zero-shot settings. We further fine-tuned TimeDiT, motivated by both its strong zero-shot performance and the inherent challenges in probabilistic modeling, where models typically struggle with calibration issues—either producing overconfident predictions with overly narrow distributions or underconfident predictions with excessively wide intervals due to insufficient domain-specific knowledge. The CRPS_sum evaluation reveals that fine-tuned TimeDiT surpasses specialized probabilistic forecasting models across multiple datasets, demonstrating exceptional performance in both zero-shot generalization and fine-tuned specialization. These results establish TimeDiT as a significant advancement in time series modeling, particularly for applications requiring robust uncertainty quantification.
Table 3: Imputation results on multivariate time series averaged over the four mask ratios.
Metrics Datasets ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity
MSE PatchTST 0.115 0.065 0.047 0.029 0.034 0.072
TimesNet 0.078 0.049 0.027 0.022 0.030 0.092
GPT4TS 0.069 0.048 0.028 0.021 0.031 0.090
Timer 0.145 0.077 0.051 0.035 0.108 0.097
TimeMixer 0.119 0.064 0.051 0.028 0.031 0.061
iTransformer 0.149 0.150 0.071 0.083 0.053 0.099
TimeDiT 0.036 0.031 0.022 0.034 0.031 0.068
MAE PatchTST 0.224 0.163 0.140 0.102 0.055 0.183
TimesNet 0.187 0.146 0.107 0.088 0.054 0.210
GPT4TS 0.173 0.141 0.105 0.084 0.056 0.207
Timer 0.243 0.172 0.141 0.105 0.168 0.194
TimeMixer 0.226 0.157 0.143 0.093 0.049 0.164
iTransformer 0.270 0.271 0.185 0.192 0.116 0.224
TimeDiT 0.122 0.111 0.093 0.104 0.036 0.172
Imputation Task:
Leveraging our model’s strong temporal understanding from forecasting tasks, rather than training a separate model or modifying the architecture, we strategically opt to fine-tune the same pre-trained checkpoint used in our forecasting experiments, maintaining architectural consistency while adapting the model’s capabilities to the imputation task. We conduct experiments on six benchmark time-series datasets: ETTh1, ETTh2, ETTm1, ETTm2, Electricity, and Weather. We use random mask ratios $\{12.5\%,25\%,37.5\%,50\%\}$ following previous studies’ settings with sequence length set to 96. Table 3 shows the imputation result averaged over the four mask ratios. TimeDiT demonstrates superior performance, achieving the best results in 10 out of 12 evaluations, while all other baselines combined secured only 2 top positions. Notably, TimeDiT achieved a 39% reduction in MSE and 22% reduction in MAE compared to the strongest baseline on the ETTh1 dataset. For full result on each mask ratio, please refer to section E.2.
Figure 2: TimeDiT with textual information.
Multimodal TimeDiT:
While textual information is intuitively crucial for precise time series analysis, effectively aligning textual and numerical data has remained challenging. To address this, we explore the integration of textual information as classifiers in TimeDiT, incorporating two key elements as guidance ($c$ in Figure 1): TSD’s frequency (Fre) for capturing temporal periodicity, and TSD’s categories (Cat) for representing domain-specific features. We pre-train three variants of TimeDiT and apply them in a zero-shot setting on Solar and Traffic datasets. The results in Figure 2 demonstrate that utilizing both types of information significantly boosts zero-shot performance, indicating TimeDiT’s capacity to leverage external information for rapid adaptation to both learned and specific representations. Comparing single-term guidance with the combined TimeDiT+Fre+Cat model reveals that precise, multi-faceted information is necessary to achieve optimal results. These experiments highlight that TimeDiT’s integration of textual context improves forecasting accuracy, enabling more informed decision-making in real-world time series applications.
Table 4: Physics-informed TimeDiT results for PDE forecasting, including both mean error and error bars. Lower values indicate better performance and closer adherence to physical laws.
@@ 4.1 Zero-Shot and Finetuning Setting
4.2 Physics-Informed TimeDiT
@@
Our approach enables the direct incorporation of physics knowledge into the pre-trained foundation model without fine-tuning. In this section, we evaluate how effectively our pre-trained foundation model can integrate physics-informed knowledge into time series forecasting. We study six 1D partial differential equations (PDEs) forecasting from  (Takamoto et al., 2022): general Navier-Stokes Equations, Kolmogorov Flow (a specific case of Navier-Stokes Equations), Advection Equations, Burgers Equations, Diffusion Soeption and Computational Fluid Dynamics (CFD). Table 4 clearly demonstrates that our proposed model editing solution, which incorporates physics knowledge, significantly outperforms previous sampling strategies introduced in DDPM (Ho et al., 2020), DDIM (Song et al., 2021), and TS Diffusion, which proposes the Self-Guidance (Kollovieh et al., 2023) to improve sampling quality. By leveraging domain-specific physical information, our approach achieves substantial performance improvements over these baselines, highlighting the effectiveness of integrating physics-informed priors into the diffusion model sampling process. This represents a novel advance in scientific machine learning, enabling rapid adaptation to specific physical systems.
4.3 In Domain Setting
@@
Table 5: Anomaly Detection result on 100-length multivariate time series. We calculate F1 score as % for each dataset. ’.’ notation in model name stands for transformer.
Methods TimeDiT TimeMixer iTrans. GPT4TS TimesNet PatchTS.
MSL 89.33 81.95 72.54 82.45 81.84 78.70
SMAP 95.91 67.63 66.76 72.88 69.39 68.82
SWaT 96.46 88.84 92.63 94.23 93.02 85.72
SMD 83.28 78.33 82.08 86.89 84.61 84.62
PSM 97.57 93.11 95.32 97.13 97.34 96.08
1st Pl Count 4 0 0 1 0 0
Anomaly Detection Task:
We conduct experiments on five real-world datasets from industrial applications: MSL, SMAP, SWaT, SMD, and PSM. Given the distinct characteristics and patterns of anomalies in different industrial systems, in-domain evaluation is crucial for assessing model robustness and reliability in practical settings. As shown in Table 5, TimeDiT outperforms baseline models on four of the five datasets. Notably, on the SMAP dataset, TimeDiT achieves a remarkable 23.03-point improvement in F1 score compared to the previous best baseline. These results demonstrate the effectiveness of our approach in handling real-world anomaly detection scenarios across various industrial applications.
Table 6: Synthetic Generation results on 24-length multivariate time series. We calculate discriminative and predictive scores according to (Yoon et al., 2019).
@@
Table 7: Forecasting results on practical scenarios with both deterministic metric (MAE/MSE) for accuracy evaluation and probabilistic metric (CRPS/CRPS_sum) for uncertainty quantification. Bold indicates best result, Underline indicates the second best result.
Air Quality MIMIC-III PhysioNet(a) PhysioNet(b) PhysioNet(c) NASDAQ
MAE/MSE MAE/MSE MAE/MSE MAE/MSE MAE/MSE MAE/MSE
DLinear 0.683/0.685 0.786/1.000 0.686/0.758 0.733/0.922 0.715/0.813 2.715/8.137
Neural ODE 0.678/0.679 0.784/0.999 0.685/0.756 0.732/0.918 0.713/0.811 3.227/11.155
Neural CDE 0.683/0.685 0.787/1.002 0.688/0.754 0.733/0.921 0.713/0.814 3.319/11.816
PatchTST 0.685/0.683 0.778/0.987 0.699/0.780 0.733/0.932 0.714/0.802 3.182/10.635
GPT4TS 0.696/0.701 0.750/0.921 0.697/0.772 0.734/0.921 0.713/0.817 3.176/10.873
CSDI 0.539/0.554 0.551/0.681 0.548/0.548 0.665/0.792 0.665/0.695 0.524/0.388
DiffTS 0.521/0.538 0.677/0.908 0.610/0.742 0.701/0.880 0.678/0.872 1.951/9.515
TimeMixer 0.691/0.697 0.769/0.981 0.692/0.775 0.734/0.920 0.707/0.805 3.267/11.511
TimeLLM 0.701/0.705 0.787/1.020 0.687/0.761 0.731/0.931 0.713/0.800 3.125/10.276
MG-TSD 0.471/0.364 - - - - 0.522/3.324
TimeDiT 0.457/0.354 0.517/0.534 0.577/0.620 0.659/0.766 0.543/0.561 0.516/0.418
CRPS/_sum CRPS/_sum CRPS/_sum CRPS/_sum CRPS/_sum CRPS
DLinear 0.662/0.544 0.770/0.748 0.764/0.812 0.794/0.793 0.767/0.797 0.342
Neural ODE 0.657/0.529 0.769/0.733 0.763/0.806 0.792/0.789 0.765/0.793 0.426
Neural CDE 0.659/0.551 0.771/0.754 0.763/0.799 0.792/0.786 0.765/0.791 0.439
PatchTST 0.664/0.564 0.771/0.721 0.769/0.812 0.791/0.775 0.766/0.777 0.410
GPT4TS 0.666/0.584 0.751/0.690 0.767/0.809 0.795/0.798 0.770/0.768 0.419
CSDI 0.598/0.620 0.504/0.798 0.620/0.641 0.725/0.787 0.669/0.748 0.096
DiffTS 0.649/0.719 0.633/0.676 0.628/0.668 0.720/0.724 0.679/0.719 0.283
TimeMixer 0.667/0.576 0.776/0.724 0.763/0.805 0.794/0.798 0.757/0.784 0.432
TimeLLM 0.664/0.571 0.785/0.700 0.752/0.797 0.795/0.795 0.757/0.754 0.405
MG-TSD 0.579/0.564 - - - - 0.275
TimeDiT 0.554/0.522 0.599/0.649 0.616/0.640 0.708/0.710 0.668/0.708 0.091
Synthetic Generation Task:
@@
We conduct experiments to synthesize multivariate time series and evaluate performance using the discriminative score and predictive score metrics under a "train on synthetic test on real" experimental setup with sequence length set to 24  (Yuan and Qiao, 2024). Table 6 shows the result on synthetic generation where TimeDiT, in general, consistently generates more realistic synthetic samples compared to baselines, even on challenging energy datasets. This demonstrates TimeDiT’s strength in complex time series synthesis. PCA visualization of synthesis performance in Appendix E.3 shows that TimeDiT’s samples markedly overlap the original data distribution better than other methods. Qualitative and quantitative results confirm TimeDiT’s superior ability to model intricate characteristics for realistic time series synthesis, even on multidimensional, complex datasets.
Practical Scenarios: Missing Data and Multi-Resolution Forecasting.
To evaluate TimeDiT’s performance in realistic scenarios, we conducted experiments incorporating three real-world challenges: missing values (validated on Air Quality, MIMIC), irregularly sampled time series (Jeon et al., 2022; Naiman et al., 2024) with varying time intervals between observations (evaluated on PhysioNet), multi-resolution data (tested on NASDAQ). This domain-specific testing ensures TimeDiT can handle both technical challenges (missing values, irregular sampling, multi-resolution) and domain-specific nuances, validating its practical utility across diverse applications. We evaluated forecasting accuracy using Mean Absolute Error (MAE) and Mean Squared Error (MSE), while uncertainty quantification (UQ) was assessed using Continuous Ranked Probability Score (CRPS) and CRPS_sum. Results in Table LABEL:tab:1_MS_RES demonstrate that TimeDiT not only achieves high accuracy in point forecasts but also provides well-calibrated probabilistic forecasts, effectively capturing the inherent uncertainties in complex time series data. The model’s strong performance in probabilistic metrics indicates its ability to generate reliable prediction intervals and accurately represent the full predictive distribution. This robust UQ capability, coupled with TimeDiT’s ability to handle missing values and irregular samples without additional designs for interpolation, positions it as a powerful tool for decision-making in uncertain environments.
4.4 Ablation Study
Table 8: Ablation study on zero-shot forecasting
Dataset TimeDiT w/o Random M. w/o Stride M. w/o Block M. w/o Phys
Solar 0.424 0.465 0.469 0.862 0.445
Electricity 0.030 0.035 0.037 0.101 0.033
Dataset Dual-attention Channel-wise Patch Token Additive Cross-attention
Solar 0.467 0.461 0.874 0.677 0.711
Electricity 0.037 0.039 0.145 0.079 0.077
As shown in Table 8, we conduct comprehensive ablation studies to validate the effectiveness of TimeDiT’s key components. The full TimeDiT model achieves superior performance (MSE of 0.424 for Solar and 0.030 for Electricity datasets) compared to its variants. Among the masking strategies, stride masking proves to be the most crucial component, as its removal leads to substantial performance degradation (MSE increasing to 0.862 for Solar and 0.101 for Electricity). In the attention mechanism analysis, our proposed temporal-attention architecture demonstrates consistent advantages over alternative designs, with patch token attention yielding the least favorable results (MSE of 0.874 and 0.145 for Solar and Electricity, respectively). Furthermore, the conditional AdaLN design emerges as the optimal choice among various conditioning approaches, reinforcing our architectural decisions.